# Non-transitive pseudo-Anosov flows

### v1

@@ 7. Action at infinity determines the plane
A consequence of the main theorem of [BFM22] is that the transitive pseudo-Anosov flows on a compact 3-manifold $M$ are completely classified up to orbit equivalence by the induced actions of $\pi_{1}(M)$ on the circles at infinity of their orbit spaces ([BFM22, Theorem 1.5]). In this section we discuss the boundary action in the general (possibly non-transitive) case for Anosov-like actions and give a simpler, unified proof that the action at infinity determines the flow in all settings.
We will not recall all the details of the construction and properties of $\partial P$ here, but refer the reader to Section 3 of [BFM22] and references therein. In particular, we will use the following result. Although proved for transitive Anosov-like actions in [BFM22], the proof goes through with only the axioms we assume here.
Proposition 7.1 (See [BFM22] Proposition 3.6).
@@
    If $g$ fixes a non-corner point $x$, then the only points of $\partial P$ fixed by $g$ are endpoints of ±$\mathcal{F}^{\pm}(x)$.
@@
The argument for the above proposition appears within the proof of Theorem 6.1 of [BFM22], we outline it here for completeness and refer the reader to [BFM22] for details.
@@
By contradiction, assume that the orbit under $G$ of $\xi$ is not dense in $\partial P$. Pick $I$ an open interval in the complement of $\overline{G\xi}$. By density of non-corner fixed leaves $\mathcal{F}^{+}$-leaves (Corollary 6.8), we can find two distinct non corner points $x,y$ fixed by respectively $g,h$ and such that $\mathcal{F}^{+}(x)$ and $\mathcal{F}^{+}(y)$ have an ideal point in $I$. By definition of $I$, its two endpoints must be fixed by both $g$ and $h$. But, since $x,y$ are nonsingular, non-corners, the only points on $\partial P$ fixed by $g$ and $h$ are the four endpoints of ±$\mathcal{F}^{\pm}(x)$ and ±$\mathcal{F}^{\pm}(y)$ respectively (Proposition 7.1), and these endpoints must be distinct (otherwise, one of ±$\mathcal{F}^{\pm}(x)$ would have to make a perfect fit with another leaf, contradicting the fact that $x$ is non-corner). ∎
Another consequence of the density of simple leaves, will be that Anosov-like actions which induces conjugated actions on the circle at infinity are actually conjugated. In order to get this result, we need the following theorem obtained in [BBM24].
Theorem 7.5 (Corollary E of [BBM24]).
@@
Note that in the special case of transitive Anosov-like actions, the proof we obtain here is much easier than the one given in [BFM22], since we do not use the main result of [BFM22]. Moreover, density of non-corner fixed leaves is much simpler to prove for transitive actions (compare [BFM22, Lemma 2.30] with Proposition 6.5 here).
@@ 8.2. Toolkit: building pseudo-Anosov flows from fatgraphs
For the next class of examples, we will use a technique developed in [BF13] to produce examples of pseudo-Anosov flows with periodic Seifert pieces by gluing building blocks together using the combinatorial data of a fatgraph. See also [BF15, BFM23].
@@
Proposition 8.6 (Barbot–Fenley [BF13], see also [BFM23]).
@@
Lemma 5.2 of [BFM23] shows that (up to an isotopy of $h$) the flow $(M,\phi)$ obtained from $(M_{i},\phi_{i})$ by gluing the boundaries according to $h$ is hyperbolic on its invariant set aside from the finitely many prong orbits corresponding to the prong orbits of the $\phi_{i}$.
@@ 8.3. Pseudo-Anosov flows, Smale classes and Smale order
Finally, we also give an example which is chain-recurrent but non-transitive.
@@
Figure 15. Fatgraph with $4$ vertices on a $8$-punctured torus (not all the edges are drawn)
@@ 8.4. An action with a totally ideal quadrilateral
In fact, if $M_{1},M_{2}$ are compact 3-manifolds with $-\mathbb{R}$ covered Anosov flows, and $G_{i}=\pi_{1}(M_{i})$, then there exists a bifoliated plane $P$ and an action of $G_{1}\ast G_{2}$ on $P$ satisfying (A1)(A4) with wandering totally ideal quadrilaterals.
@@
Let $(P,\mathcal{F}^{+},\mathcal{F}^{-})$ be a bifoliated plane and $S$ some collection of leaves of $\mathcal{F}^{+}$. We say that another plane $(\hat{P},\hat{\mathcal{F}}^{+},\hat{\mathcal{F}}^{-})$ is a blowup of $P$ along $S$ if there exists a continuous map $h\colon\hat{P}\to P$ with the following properties
@@
  • For each $l\in S$, $h^{-1}(l)$ is homeomorphic to $[0,1]\times\mathbb{R}$ via a homeomorphism sending leaves of $\hat{\mathcal{F}}^{+}$ to the vertical foliation and leaves of $\hat{\mathcal{F}}^{-}$ to the horizontal.
  • using the coordinates above, $h$ is constant on each $\hat{\mathcal{F}}^{-}$ (horizontal) leaf $[0,1]\times\{x\}$ in each $h^{-1}(l)$.
@@
Construction 8.13 (Blow-up with boundary).
Let $\hat{P}$ be a blow-up along $G\cdot l_{0}$ as above. The blow-up with boundary is a bifoliated space obtained as follows. For each $l\in G\cdot l_{0}$, we have that $h^{-1}(l)\cong\mathbb{R}\times[0,1]$. We would like to add a leaf of $\mathcal{F}^{-}$ at +$\infty$+ ∞ and at $\infty$- ∞ in the $\mathbb{R}$-coordinate to each such region so the result is homeomorphic to a totally ideal quadrilateral. To do this, one can put a metric on $\hat{P}$ so that the preimages of leaves $l\in Gl_{0}$ under $h$ (each homeomorphic to $[0,1]\times\mathbb{R}$) have the length of leaf $[0,1]\times\{x\}$ approaching $\infty$ as $x\to\pm\infty$, and then compactify this strip by adding two leaves homeomorphic to $\mathbb{R}\times\{+\infty\}$ and ${}\mathbb{R}\times\{-\infty\}$, one at either end. We call the result a Blow-up of $P$ with boundary. This is no longer a topological plane, but is homeomorphic to a simply connected subset of the plane.
@@
Start with an Anosov-like action of a group $G$ on a bifoliated plane $(P,\mathcal{F}^{+},\mathcal{F}^{-})$ coming from the orbit space of a non $\mathbb{R}$-covered Anosov flow, and fix a leaf $l_{0}$ with trivial stabilizer. Let $\bar{P}$ be the blow-up with boundary along $G\cdot l_{0}$ as in Construction 8.13. Take another bifoliated plane $(P^{\prime},\mathcal{F}^{\prime+},\mathcal{F}^{\prime-})$, and cut $P^{\prime}$ along a leaf $l^{-}$ of $\mathcal{F}^{\prime-}$, obtaining two half-planes $H_{+},H_{-}$. For each $g\in G$, take copies $H_{\pm}(g)$ of $H_{+}$ and $H_{-}$ and glue them to the compactified ends of $h^{-1}(gl)$, equivariantly with respect to the action of $G$. The result is a bifoliated plane to which the action of $G$ extends and each compactified $h^{-1}(gl)$ is a totally ideal quadrilateral. However, this action unfortunately will not satisfy all of the required axioms – in particular, it will never satisfy Axiom (A2) since the set of $\mathcal{F}^{+}$ leaves passing through one of these ideal quadrilaterals gives an open set of leaves which are fixed by no nontrivial element.
Axiom (A1) does hold, however, thanks to the choice of $l$ with trivial stabilizer. If $P^{\prime}$ contains no singular points, axiom (A3) will also be satisfied. But (A4) is again problematic: if $l$ makes a perfect fit with another leaf $f^{-}$ in $P$, then $l^{-}$ and (the preimage under $h$ of) $f^{-}$ will be non-separated, but not fixed by any element, failing Axiom (A4). (Note that, since there are no infinite product regions, as we started from a $-\mathbb{R}$-covered flow, a copy of $l^{-}$ can only be non-separated with a leaf making a perfect fit with $l$.) To avoid this, we use the following lemma, which follows from an argument of Fenley in [Fen16].
@@
Let $o$ be a dense orbit in $M$, and let $p\in P$ be the image of a lift of $o$ to $\widetilde{M}$ projected to the orbit space. Since $o$ is dense, $G\cdot p$ is dense in $P$. We claim that ±$\mathcal{F}^{\pm}(p)$ cannot make any perfect fit.
Suppose for contradiction that one of ±$\mathcal{F}^{\pm}(p)$ makes a perfect fit with another leaf. The proof of Proposition 5.5 in [Fen16] shows that $p$ is then accumulated by corners of lozenges. We sketch this for completeness: using the closing lemma one can find a periodic orbit $\delta$ which admits a lift to $\widetilde{M}$ that projects to a point $\widetilde{\delta}$ in the orbit space (notation here is borrowed from [Fen16]), close to $p$. Recurrence (here given by density of $o$) in fact allows us to find such points arbitrarily close to $p$. This periodic orbit is associated to some $g\in G$ which fixes $\widetilde{\delta}$ and can be taken (again) to move $p$ an arbitrarily small amount. Moreover, one can specify on which side of $\mathcal{F}^{s}(p)$ the image $g\mathcal{F}^{s}(p)$ lies. Considering the arrangement of leaves in the orbit space, as in the left side of Figure 9 of [Fen16], (here $p_{1}$ is labeling the projection of $p$, and $g(p_{2})$ its image under $g$) one concludes that $\widetilde{\delta}$ makes a perfect fit, and is the corner of a lozenge. Thus, $p$ is approximated by corners of lozenges.
Now by Lemma 2.30 of [BFM22], the set of points in the orbit space of a transitive, $-\mathbb{R}$-covered Anosov flow that are corners of lozenges is nowhere dense. This contradicts the density of $G\cdot p$ in $P$. ∎
With this in hand, we can now prove the main result of this section. The general idea is as follows: first, blow up an orbit $G_{1}\cdot l_{1}$ in $P_{1}$ using the blow-up with boundary construction, and $G_{2}\cdot l_{2}$ in $P_{2}$. We then glue copies of these together along the blown-up quadrilaterals in a way so that the actions of $G_{1}$ and $G_{2}$ extend equivariantly, and play “ping-pong” on the resulting space. The proof below consists in making this idea precise. ${}^{\text{\tiny\rm 1}}$ margin: 1) K: added explanation
@@
Step 1: taking many copies of blow-ups with boundary. Let $\bar{P}_{1},\bar{\mathcal{F}}^{\pm}_{1}$ and $\bar{P}_{2},\bar{\mathcal{F}}^{\pm}_{2}$ denote the blow-ups with boundary of $P_{i}$ along $G_{i}(l_{i})$, respectively, as described in Construction 8.13. Of particular importance to us is the fact that the ¯$\bar{\mathcal{F}}_{1}^{-}$ leaf space of $\bar{P}_{1}$ is identical to that of the original $\mathcal{F}_{1}^{-}$ leaf space in $P_{1}$ itself. Similarly, each ¯$\bar{\mathcal{F}}_{2}^{+}$ leaf of $\bar{P}_{2}$ is naturally identified with a leaf of $\mathcal{F}_{2}^{+}$ in $P_{2}$.
@@
Step 2: gluing along ideal quadrilaterals. Let $Q_{i}\subset\bar{P}_{i}$, $i=1,2$, be the ideal quadrilaterals corresponding to the compactified blow-ups of $l_{i}$. Let $\Phi\colon Q_{1}\to Q_{2}$ be a homeomorphism sending leaves of ¯$\bar{\mathcal{F}}^{\pm}_{1}$ to ¯$\bar{\mathcal{F}}^{\pm}_{2}$ (preserving $+$ and --). For an image $aQ_{1}$ of $Q_{1}$ under some element $a\in G_{1}$, let $\Phi_{a}$ be the gluing map $aQ_{1}\to Q_{2}$ defined by $\Phi_{a}(x)=a\Phi a^{-1}(x)$. Similarly for $b\in G_{2}$ let $\Phi_{b}\colon Q_{1}\to bQ_{2}$ denote the homeomorphism whose inverse $\Phi_{b}^{-1}\colon bQ_{2}\to Q_{1}$ is given by $\Phi_{b}^{-1}(x)=b\Phi^{-1}b^{-1}(x)$.

### v2

@@ 7. Action at infinity determines the plane
A consequence of the main theorem of [BFM25b] is that the transitive pseudo-Anosov flows on a compact 3-manifold $M$ are completely classified up to orbit equivalence by the induced actions of $\pi_{1}(M)$ on the circles at infinity of their orbit spaces ([BFM25b, Theorem 1.5]). In this section we discuss the boundary action in the general (possibly nontransitive) case for Anosov-like actions and give a simpler, unified proof that the action at infinity determines the flow in all settings.
We will not recall all the details of the construction and properties of $\partial P$ here, but refer the reader to Section 3 of [BFM25b] and references therein. In particular, we will use the following result. Although proved for transitive Anosov-like actions in [BFM25b], the proof goes through with only the axioms we assume here.
Proposition 7.1 (See [BFM25b] Proposition 3.6).
@@
    If $g$ fixes a non-corner point $x$, then the only points of $\partial P$ fixed by $g$ are endpoints of $\mathcal{F}^{\pm}(x)$.
@@
The argument for the above proposition appears within the proof of Theorem 6.1 of [BFM25b], we outline it here for completeness and refer the reader to [BFM25b] for details.
@@
By contradiction, assume that the orbit under $G$ of $\xi$ is not dense in $\partial P$. Pick $I$ an open interval in the complement of $\overline{G\xi}$. By density of non-corner fixed leaves $\mathcal{F}^{+}$-leaves (Corollary 6.11), we can find two distinct non corner points $x,y$ fixed by respectively $g,h$ and such that $\mathcal{F}^{+}(x)$ and $\mathcal{F}^{+}(y)$ have an ideal point in $I$. By definition of $I$, its two endpoints must be fixed by both $g$ and $h$. But, since $x,y$ are nonsingular, non-corners, the only points on $\partial P$ fixed by $g$ and $h$ are the four endpoints of $\mathcal{F}^{\pm}(x)$ and $\mathcal{F}^{\pm}(y)$ respectively (Proposition 7.1), and these endpoints must be distinct (otherwise, one of $\mathcal{F}^{\pm}(x)$ would have to make a perfect fit with another leaf, contradicting the fact that $x$ is non-corner). ∎
Another consequence of the density of simple leaves, will be that Anosov-like actions which induces conjugated actions on the circle at infinity are actually conjugated. In order to get this result, we need the following theorem obtained in [BBM25].
Theorem 7.5 (Corollary E of [BBM25]).
@@
Note that in the special case of transitive Anosov-like actions, the proof we obtain here is much easier than the one given in [BFM25b], since we do not use the main result of [BFM25b]. Moreover, density of non-corner fixed leaves is much simpler to prove for transitive actions (compare [BFM25b, Lemma 2.30] with Proposition 6.8 here).
@@ 8.2. Toolkit: building pseudo-Anosov flows from fatgraphs
For the next class of examples, we will use a technique developed in [BF13] to produce examples of pseudo-Anosov flows with periodic Seifert pieces by gluing building blocks together using the combinatorial data of a fatgraph. See also [BF15, BFM25a].
@@
Proposition 8.6 (Barbot–Fenley [BF13], see also [BFM25a]).
@@
Lemma 5.2 of [BFM25a] shows that (up to an isotopy of $h$) the flow $(M,\phi)$ obtained from $(M_{i},\phi_{i})$ by gluing the boundaries according to $h$ is hyperbolic on its invariant set aside from the finitely many prong orbits corresponding to the prong orbits of the $\phi_{i}$.
@@ 8.3. Pseudo-Anosov flows, Smale classes and Smale order
Finally, we also give an example which is chain-recurrent but nontransitive.
@@
Figure 15. Fatgraph with 44 vertices on a 88-punctured torus (not all the edges are drawn)
Our last example involving Smale classes is the following, illustrating the importance of considering singular Smale classes.
Proposition 8.10.
There exists pseudo-Anosov flows on non-compact $3$-manifolds whose non-wandering set consists entirely of isolated prong singularities.
Proof.
The construction is similar to that of Example 8.4. Consider an admissible fat graph on a surface $S$ such that all vertices have valence >4>4. The associated flow hence has only prong singularities, and we build a pseudo-Anosov flow $\phi$ on a compact manifold $M$ by taking two copies of that associated flow and gluing them along their incoming/outgoing boundaries.
Call $k$ the number of boundary components of $S$, so that $M$ has $k$ torus in its JSJ decomposition, that we denote by $T_{1},\dots,T_{k}$. Let $G<\pi_{1}(M)$ be the kernel of the morphism $\pi_{1}(M)\to\mathbb{Z}^{k}$ whose $i$-th coordinate is given by the intersection number of $g\in\pi_{1}(M)$ with the torus $T_{i}$.
Then $G$ corresponds to the fundamental group of an infinite (regular) cover $\widehat{M}$ of $M$, and any element $g\in G$ that represents a periodic orbit must by construction corresponds to an element of the fundamental group of one of the pieces of the JSJ decomposition of $\widehat{M}$, so must represent a prong singularity associated to one of the vertices of the original fat graph on $S$. Hence, the lift $\widehat{\phi}$ has only isolated prong singularities.
As in Example 8.4 one can easily verify that the action of $G$ on the orbit space is Anosov-like. ∎
@@ 8.4. An action with a totally ideal quadrilateral
In fact, if $M_{1},M_{2}$ are compact 3-manifolds with non-$\mathbb{R}$ covered Anosov flows, and $G_{i}=\pi_{1}(M_{i})$, then there exists a bifoliated plane $P$ and an action of $G_{1}\ast G_{2}$ on $P$ satisfying (A1)(A4) with wandering totally ideal quadrilaterals.
@@
[…]
@@
  • For each $l\in S$, $h^{-1}(l)$ is homeomorphic to $0,1]\times\mathbb{R}$ via a homeomorphism sending leaves of $\hat{\mathcal{F}}^{+}$ to the vertical foliation and leaves of $\hat{\mathcal{F}}^{-}$ to the horizontal.
  • using the coordinates above, $h$ is constant on each $\hat{\mathcal{F}}^{-}$ (horizontal) leaf $0,1]\times\{x\}$ in each $h^{-1}(l)$.
@@
Construction 8.14 (Blow-up with boundary).
Let $\hat{P}$ be a blow-up along $G\cdot l_{0}$ as above. The blow-up with boundary is a bifoliated space obtained as follows. For each $l\in G\cdot l_{0}$, we have that $h^{-1}(l)\cong\mathbb{R}\times[0,1]$. We would like to add a leaf of $\mathcal{F}^{-}$ at $\infty$ and at -$\infty$ in the $\mathbb{R}$-coordinate to each such region so the result is homeomorphic to a totally ideal quadrilateral. To do this, one can put a metric on $\hat{P}$ so that the preimages of leaves $l\in Gl_{0}$ under $h$ (each homeomorphic to $0,1]\times\mathbb{R}$) have the length of leaf $0,1]\times\{x\}$ approaching $\infty$ as $x\to\pm\infty$, and then compactify this strip by adding two leaves homeomorphic to $\mathbb{R}\times\{+\infty\}$ and $\mathbb{R}\times\{-\infty\}$, one at either end. We call the result a Blow-up of $P$ with boundary. This is no longer a topological plane, but is homeomorphic to a simply connected subset of the plane.
@@
Start with an Anosov-like action of a group $G$ on a bifoliated plane $(P,\mathcal{F}^{+},\mathcal{F}^{-})$ coming from the orbit space of a non $\mathbb{R}$-covered Anosov flow, and fix a leaf $l_{0}$ with trivial stabilizer. Let $\bar{P}$ be the blow-up with boundary along $G\cdot l_{0}$ as in Construction 8.14. Take another bifoliated plane $(P^{\prime},\mathcal{F}^{\prime+},\mathcal{F}^{\prime-})$, and cut $P^{\prime}$ along a leaf $l^{-}$ of $\mathcal{F}^{\prime-}$, obtaining two half-planes $H_{+},H_{-}$. For each $g\in G$, take copies $H_{\pm}(g)$ of $H_{+}$ and $H_{-}$ and glue them to the compactified ends of $h^{-1}(gl)$, equivariantly with respect to the action of $G$. The result is a bifoliated plane to which the action of $G$ extends and each compactified $h^{-1}(gl)$ is a totally ideal quadrilateral. However, this action unfortunately will not satisfy all of the required axioms – in particular, it will never satisfy Axiom (A2) since the set of $\mathcal{F}^{+}$ leaves passing through one of these ideal quadrilaterals gives an open set of leaves which are fixed by no nontrivial element.
Axiom (A1) does hold, however, thanks to the choice of $l$ with trivial stabilizer. If $P^{\prime}$ contains no singular points, axiom (A3) will also be satisfied. But (A4) is again problematic: if $l$ makes a perfect fit with another leaf $f^{-}$ in $P$, then $l^{-}$ and (the preimage under $h$ of) $f^{-}$ will be non-separated, but not fixed by any element, failing Axiom (A4). (Note that, since there are no infinite product regions, as we started from a non-$\mathbb{R}$-covered flow, a copy of $l^{-}$ can only be non-separated with a leaf making a perfect fit with $l$.) To avoid this, we use the following lemma, which follows from an argument of Fenley in [Fen16].
@@
Let $o$ be a dense orbit in $M$, and let $p\in P$ be the image of a lift of $o$ to $\widetilde{M}$ projected to the orbit space. Since $o$ is dense, $G\cdot p$ is dense in $P$. We claim that $\mathcal{F}^{\pm}(p)$ cannot make any perfect fit.
Suppose for contradiction that one of $\mathcal{F}^{\pm}(p)$ makes a perfect fit with another leaf. The proof of Proposition 5.5 in [Fen16] shows that $p$ is then accumulated by corners of lozenges. We sketch this for completeness: using the closing lemma one can find a periodic orbit $\delta$ which admits a lift to $\widetilde{M}$ that projects to a point $\widetilde{\delta}$ in the orbit space (notation here is borrowed from [Fen16]), close to $p$. Recurrence (here given by density of $o$) in fact allows us to find such points arbitrarily close to $p$. This periodic orbit is associated to some $g\in G$ which fixes $\widetilde{\delta}$ and can be taken (again) to move $p$ an arbitrarily small amount. Moreover, one can specify on which side of $\mathcal{F}^{s}(p)$ the image $g\mathcal{F}^{s}(p)$ lies. Considering the arrangement of leaves in the orbit space, as in the left side of Figure 9 of [Fen16], (here $p_{1}$ is labeling the projection of $p$, and $g(p_{2})$ its image under $g$) one concludes that $\widetilde{\delta}$ makes a perfect fit, and is the corner of a lozenge. Thus, $p$ is approximated by corners of lozenges.
Now by Lemma 2.30 of [BFM25b], the set of points in the orbit space of a transitive, non-$\mathbb{R}$-covered Anosov flow that are corners of lozenges is nowhere dense. This contradicts the density of $G\cdot p$ in $P$. ∎
With this in hand, we can now prove the main result of this section. The general idea is as follows: first, blow up an orbit $G_{1}\cdot l_{1}$ in $P_{1}$ using the blow-up with boundary construction, and $G_{2}\cdot l_{2}$ in $P_{2}$. We then glue copies of these together along the blown-up quadrilaterals in a way so that the actions of $G_{1}$ and $G_{2}$ extend equivariantly, and play “ping-pong” on the resulting space. The proof below consists in making this idea precise.
@@
Step 1: taking many copies of blow-ups with boundary. Let $\bar{P}_{1},\bar{\mathcal{F}}^{\pm}_{1}$ and $\bar{P}_{2},\bar{\mathcal{F}}^{\pm}_{2}$ denote the blow-ups with boundary of $P_{i}$ along $G_{i}(l_{i})$, respectively, as described in Construction 8.14. Of particular importance to us is the fact that the $\bar{\mathcal{F}}_{1}^{-}$ leaf space of $\bar{P}_{1}$ is identical to that of the original $\mathcal{F}_{1}^{-}$ leaf space in $P_{1}$ itself. Similarly, each $\bar{\mathcal{F}}_{2}^{+}$ leaf of $\bar{P}_{2}$ is naturally identified with a leaf of $\mathcal{F}_{2}^{+}$ in $P_{2}$.
@@
Step 2: gluing along ideal quadrilaterals. Let $Q_{i}\subset\bar{P}_{i}$, $i=1,2$, be the ideal quadrilaterals corresponding to the compactified blow-ups of $l_{i}$. Let $\Phi\colon Q_{1}\to Q_{2}$ be a homeomorphism sending leaves of $\bar{\mathcal{F}}^{\pm}_{1}$ to $\bar{\mathcal{F}}^{\pm}_{2}$ (preserving ++ and -). For an image $aQ_{1}$ of $Q_{1}$ under some element $a\in G_{1}$, let $\Phi_{a}$ be the gluing map $aQ_{1}\to Q_{2}$ defined by $\Phi_{a}(x)=a\Phi a^{-1}(x)$. Similarly for $b\in G_{2}$ let $\Phi_{b}\colon Q_{1}\to bQ_{2}$ denote the homeomorphism whose inverse $\Phi_{b}^{-1}\colon bQ_{2}\to Q_{1}$ is given by $\Phi_{b}^{-1}(x)=b\Phi^{-1}b^{-1}(x)$.